OpenStudy (anonymous):

what would the table graph look like for f(x)=(3/4)x ?

OpenStudy (anonymous):

@ivettef365 would u happen to know this ?

OpenStudy (anonymous):

what curve would it make..

OpenStudy (anonymous):

write down a table, remember the formula y=mx+b, here you don't have a b, and m=slope, so i.e. if x=1, then y=3/4, these are your points on the graph which you will then trace to see the graph

OpenStudy (anonymous):

would it be a up slope or dwn?

OpenStudy (anonymous):

m is a positive, therefore the slope is..?

OpenStudy (anonymous):

would look just like that correct?

OpenStudy (anonymous):

your question is fro f(x)=(3/4)x, right

OpenStudy (anonymous):